OpenStudy (openstudygirl2):

Describe a real-world situation in which opposite quantities combine to make zero.

OpenStudy (openstudygirl2):

Help me @agent0smith

OpenStudy (skullpatrol):

How about a debt of $50 combined with a gift of $50?

OpenStudy (openstudygirl2):

Yeah maybe!

OpenStudy (skullpatrol):

You could walk south for 5 miles and then turn around and walk north for 5 miles.

OpenStudy (openstudygirl2):

Can i ask another question ?

OpenStudy (skullpatrol):

Sure :)

OpenStudy (openstudygirl2):

Describe a recent situation where you added or subtracted.

OpenStudy (skullpatrol):

I checked the grocery bill by adding up the items.

OpenStudy (skullpatrol):

Any deposit or withdrawal from my bank account better add up.

OpenStudy (openstudygirl2):

OKay

OpenStudy (openstudygirl2):

How is the additive inverse property used to subtract rational numbers?

OpenStudy (skullpatrol):

Do you have any ideas?

OpenStudy (openstudygirl2):

Not a Clue! xD

OpenStudy (skullpatrol):

Recall the additive inverse property $$\large a+(-a)=0$$

OpenStudy (openstudygirl2):

Okay

OpenStudy (skullpatrol):

$$\large a+(-a)=a-a=0$$

OpenStudy (skullpatrol):

That^ tells you to subtract a number you add its additive inverse.

OpenStudy (openstudygirl2):

Okay! But what would i write for my answer ?

OpenStudy (skullpatrol):

$$\large\dfrac{a}{c}-\dfrac{b}{c}=\dfrac{a-b}{c}=\dfrac{a+(-b)}{c}$$

OpenStudy (skullpatrol):

The last step a + (-b) uses the additive inverse of b which is -b

OpenStudy (skullpatrol):

To subtract b add the additive inverse of b.

OpenStudy (openstudygirl2):

Okay :D